OpenStudy (anonymous):

sin(-420) * cos(390) +cos(-660)* sin(330) is ?

OpenStudy (mani_jha):

-sin(360+60)cos(360+30)+cos(2*360-60)sin(360-30) =-sin60cos30+cos60(-sin30) =-sin60cos30-sin30cos60 =-sin(60+30) =-sin90 =-1 ok?

OpenStudy (anonymous):

Thanzz